OpenStudy (anonymous):

Find the 8th term of the geometric sequence show below −4x2,12x6,−36x10,... Common Ratio: r=a2a1 r=12x6−4x2=−3x4 an=a1⋅rn−1 General Formula a8=−4x2(−3x4)7 a8=−4x2(−2187x28) Exponent a8=8748x30 I do not understand how you get -2187

OpenStudy (anonymous):

The numbers after the x's are exponents

OpenStudy (shaik0124):

r=a2/a1 =72/-8 =-9

OpenStudy (shaik0124):

a8=-8*(-9)^7

OpenStudy (shaik0124):

solve this u will get the answer

OpenStudy (shaik0124):

@gumstopper getting or not??